OpenStudy (anonymous):

A fish swimming in a horizontal plane has velocity varrowboldi = (4.00 ihatbold + 1.00 jhatbold) m/s at a point in the ocean where the position relative to a certain rock is rarrowboldi = (16.0 ihatbold − 1.60 jhatbold) m. After the fish swims with constant acceleration for 15.0 s, its velocity is varrowbold = (25.0 ihatbold − 1.00 jhatbold) m/s. (a) What are the components of the acceleration of the fish?

OpenStudy (anonymous):

solve it component-wise: < (4.00 + a*15 ) i + (1.00 + a*15) j> = < 25 i + -1 j> i component of acceleration: 4.00 +a*15 = 25 j component of acceleration: 1.00 +a*15 = -1

OpenStudy (anonymous):

I did figure that out eventually.

OpenStudy (anonymous):

... could have closed it...

OpenStudy (anonymous):

I'm sorry...you had already posted it before I figured it out, otherwise I would have closed it.

OpenStudy (anonymous):

no problem.